OpenStudy (anonymous):

Find the average rate of change of f(x)= 2x^2 - 9x from x=2 to x=6 . Simplify your answer as much as possible.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

we need some numbers \[f(6),f(2)\] to start

OpenStudy (anonymous):

\[f(6)=2\times 6^2-9\times 6=18\] \[f(2)=2\times 2^2-9\times 2=-10\]

OpenStudy (anonymous):

what you want to compute is the slope \[\frac{f(6)-f(2)}{6-4}\] which is in this case \[\frac{18-(-10)}{6-2}\]

OpenStudy (anonymous):

i get \[\frac{28}{4}=7\]

OpenStudy (anonymous):

oh wow that makes so much more sense the explanations I had been looking up! thank you so much! that really wasn't as complicated as I was thinking

OpenStudy (anonymous):

lol no it is just the slope of the line between the points \((2,-10)\) and \((6,18)\)

OpenStudy (anonymous):

Lol everything I looked up or watched looked really drawn out =p